A particular brand of coffee contains an average of 112 mg of caffeine per cup with a standard deviation of 20 mg. A barista wants to investigate the same to estimate the true population mean caffeine content correct to within 5 mg adopting 95% confidence. How many cups of the same brand of coffee does he need for a sample?


1
Expert's answer
2021-06-22T05:57:33-0400

The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96"

The corresponding confidence interval is computed as shown below:


"CI=(x-z_c\times \dfrac{\sigma}{\sqrt{n}}, x+z_c\times \dfrac{\sigma}{\sqrt{n}})"

"SE=z_c\times \dfrac{\sigma}{\sqrt{n}}=>n=( \dfrac{z_c\sigma}{SE})^2"

Substitute


"n=( \dfrac{1.96(20)}{5})^2=61"

He needs 61 cups of the same brand of coffee for a sample.